OpenStudy (anonymous):

Solve log x = 3. x = 3 x = 30 x = 1,000 x = 10,000

OpenStudy (anonymous):

helllp

OpenStudy (anonymous):

ILL GIVE MEDALS

OpenStudy (anonymous):

@Chop-GTA5

OpenStudy (anonymous):

@geerky42

TheSmartOne (thesmartone):

If \(\Huge log _{b}y=x\) then \(\Huge y=b^x\)

TheSmartOne (thesmartone):

In your case it is \(\Huge log _{10}x=3\) So then \(\Huge x=10^3\)

TheSmartOne (thesmartone):

And \(\Huge 10^3=1000\)

TheSmartOne (thesmartone):

So your answer is \(\Huge x=1000\)